Calculate Economic Value Added (EVA) if the Earnings before Interest and tax is Rs 20, 00,000. The capital structure of the firm consists of Debt-Equity ratio of 2: 3, pretax cost of debt is 14% and the tax rate is 35%. The equity beta is 1.5. The risk free rate of return is 9% and risk premium is 5%. Total borrowed capital of the firm is Rs 45, 00,000. Comment on the value of EVA


1
Expert's answer
2021-08-05T05:52:02-0400

Economic value added (EVA) is a financial management concept that evaluates an organization's economic performance using variables such as net income, cost of capital, working capital, and taxes. EVA is determined by dividing net income by the cost of capital, the amount of working capital generated, and taxes.

EVA=EBIT(1−T)−KwC

Kw= 45000/2 *5= 112500

112500*1.6= 181750

When a corporation creates value, this is highlighted directly by this positive figure.